# Can Frontier LLMs Replace Annotators in Biomedical Text Mining? Analyzing Challenges and Exploring Solutions

### v1

@@ 4.2 Evaluation of Reasoning Model (o1-mini)
Table 6 presents our experimental results on the reasoning model (OpenAI o1-mini). Except for the fact that the model’s built-in test-time computation replaces the need for Two-Step Inference, all prompting settings remain the same as those used with GPT-4o. Across all datasets, o1-mini performs slightly worse than GPT-4o enhanced with our proposed approach. We speculate that larger models, such as o1 or Deepseek-R1[72], which are capable of deeper reasoning, could achieve higher performance than o1-mini. However, due to computational costs and API providers’ constraints on processing time, we were unable to include these models in our experiments.
@@ 4.3 Result of Model Distillation
We followed the same procedures as outlined in the original studies of HoC and BC5CDR-RE datasets to collect papers published after April 2024, and then constructed synthetic training datasets entirely annotated by LLM. These datasets were used to fine-tune the BioLinkBERT-Large model, which was subsequently evaluated on a manually annotated gold-standard test set. As shown in Table 7 , the distilled BioLinkBERT model achieved a Macro-F1 score of 0.817 on the HoC dataset, slightly behind the performance of 0.873 when fine-tuned on the official training set which is smaller but annotated by human experts. For the BC5CDR-RE dataset, the distilled BioLinkBERT model achieved a Macro-F1 score of 0.646, closely approaching the performance of 0.650 obtained by fine-tuning on the official training set. Notably, it outperformed the fine-tuned BioGPT, which achieved a Macro-F1 score of 0.462 and was reported as the SOTA method on this dataset.
@@ 5 Conclusions and Future Work
In our study, we first analyzed the factors contributing to the suboptimal performance of LLMs when applied to biomedical text mining tasks using conventional approaches. Building on this analysis, we developed a composite prompting strategy specifically targeted to address these challenges, and our results demonstrate that the frontier LLMs like GPT-4o could challenge the prevailing notion that they are inherently unsuitable for discriminative tasks in biomedical domain. While core ideas of these prompt engineering techniques and information retrieval frameworks were not originally ours, we customized them to better suit the specific needs of the field.
Our results indicate that combining two-step inference and instruction retrieval from annotation guidelines can greatly enhance LLM performance. However, this method relies heavily on predefined rules for segmenting the annotation guidelines and extracting relevant information during inference, demanding significant preparatory and coding efforts. Future improvements could include developing an LLM annotator agent that automatically queries the necessary guidelines and examples, or integrating a trainable retriever like in Li et al.’s research[73] to dynamically extract guidelines, instead of the rule-based approach. Additionally, we did not fine-tune or deploy any open-source LLM due to computational limitation. Nonetheless, the proposed prompt engineering strategies and model fine-tuning approaches are complementary, and their combination may potentially yield even better performance.
@@
Additionally, since our evaluations were conducted on commonly used biological text databases, we cannot entirely rule out the risk of data contamination (Appendix 7.3). Considering this, we recommend continuous monitoring and management of LLM performance when deploying this methodology in production environments.
@@ 6 Data Availability
The data and prompt used in this study can be found at the following repository: https://github.com/ekkkkki/LLM-Replace-Annotators-in-Biomedical-Text-Mining.

### v2

@@ 4.2 Evaluation of Reasoning Model (o1-mini)
Table 6 presents our experimental results on the reasoning model (OpenAI o1-mini). Except for the fact that the model’s built-in test-time computation replaces the need for Two-Step Inference, all prompting settings remain the same as those used with GPT-4o. Across all datasets, o1-mini performs slightly worse than GPT-4o enhanced with proposed approach. We speculate that larger models, such as o1 or Deepseek-R1[72], which are capable of deeper reasoning, could achieve higher performance than o1-mini. However, due to computational costs and API providers’ constraints on processing time, we were unable to include these models in our experiments.
@@ 4.3 Result of Model Distillation
Following original procedures, we retrieved recent papers and constructed synthetic training datasets for HoC and BC5CDR-RE annotated by GPT-4o.
BioLinkBERT-Large fine-tuned on these synthetic datasets achieved a Macro-F1 of 0.817 on HoC, slightly below its performance on the smaller, expert-annotated official training set (Table 7). For BC5CDR-RE, the distilled BioLinkBERT model achieved a Macro-F1 of 0.646, closely approaching the performance of 0.650 obtained by fine-tuning on the official training set. Notably, it outperformed the fine-tuned BioGPT, which achieved a Macro-F1 score of 0.462 and was reported as the SOTA method on this dataset.
@@ 5 Conclusions and Future Work
Our study analyzed factors contributing to LLMs’ suboptimal performance in biomedical text mining using conventional methods. We subsequently developed targeted prompting strategies to address these challenges. While core ideas of these prompt engineering techniques and information retrieval frameworks were not originally ours, we customized them to better suit the specific needs of those tasks. Results demonstrated that frontier LLMs, like GPT-4o, could challenge the prior notion of their unsuitability for biomedical discriminative tasks.
Our results indicate that combining two-step inference and instruction retrieval from annotation guidelines can greatly enhance LLM performance. However, the current method relies heavily on predefined rules for guideline segmentation and information retrieval, requiring substantial preparatory and coding efforts. Future improvements could involve developing an LLM annotator agent that automatically queries the necessary guidelines and examples, or integrating a trainable retriever like in Li et al.’s research[73] to dynamically extract guidelines, instead of the rule-based approach. Additionally, we did not fine-tune any open-source LLM due to computational limitation, but the proposed strategies and model fine-tuning approaches are complementary, and their combination may potentially yield even better performance.
@@
Additionally, since our evaluations were conducted on commonly used biomedical text mining datasets, we cannot entirely rule out the risk of data contamination (Appendix 1.4). Considering this, we recommend continuous monitoring and management of LLM performance when deploying this methodology in production environments.
@@ 6 Data Availability
The data and prompt can be found at the following repository: https://github.com/ekkkkki/LLM-Replace-Annotators-in-Biomedical-Text-Mining.